NSU Manufacturing, Inc. produces textbooks and is preparing budgets for next year. The following information relates to next year’s activities:

Expected sales for next year are 195,000 textbooks at an estimated production cost of $32.00 per textbook and an estimated selling price of $40.00 per textbook.

Beginning inventories:

Raw materials:

Paper 25,000 yards @ $2.99 per yard

Plastic 8,000 pounds @ $5.49 per pound

Work-in-process 928 textbooks @ $28.93 per textbook

Finished goods 2,500 textbooks @ $32.00 per textbook

Desired ending inventories:

Raw materials:

Paper 30,000 yards @ $2.99 per yard

Plastic 11,000 pounds @ $5.49 per pound

Work-in-process 762 textbooks @ 27.29 per textbook

Finished goods 25,000 textbooks @ $32.00 per textbook

Direct materials needed:

Paper 3.5 yards per textbook

Plastic 0.6 pounds per textbook

Direct labor time needed: 0.33 hours per textbook @ $18.50 per direct labor hour

Variable manufacturing overheads:

Indirect materials and supplies $0.33 per textbook

Materials handling $0.45 per textbook

Other indirect labor $0.12 per textbook

Fixed manufacturing overheads:

Supervisory labor $98,000

Maintenance and repairs $53,000

Plant administration $72,000

Utilities $36,000

Depreciation $128,000

Insurance $30,000

Property taxes $54,000

Other $13,000

Variable marketing costs:

Sales commission $2.50 per textbook

Other marketing $0.89 per textbook

Fixed marketing costs:

Sales salaries $150,000

Advertising $253,000

Other $98,000

Administrative costs:

Administrative salaries $220,000

Legal and accounting staff $80,000

Data processing services $65,000

Outside professional services $52,000

Depreciation - building, furniture, and equipment $49,000

Interest expense $6,000

Property taxes $76,000

Other $24,000


Income tax rate: 35%

Find out the following:

  • Break-even units

  • Break-even price per textbook

  • Profit (loss) after tax when budgeted sales price per unit is $30.00 per textbook and budgeted sales volume is 250,000 textbooks

  • Profit (loss) after tax when budgeted sales price per unit is $35.00 per textbook and budgeted sales volume is 150,000 textbooks
Answer #1

1) Breakeven is the point where contribution from sale of products is equal to fixed costs.

Now, contribution refers to, Sales price (less) variable costs.

Calculation of fixed costs:

Particulars Amount
Fixed Manufacturing Overhead
Supervisory Labor 98000
Maintenance and Repair 53000
Plant Administration 72000
Utilities 36000
Depreciation 128000
Insurance 30000
Property Tax 54000
Others 13000
Fixed Marketing Costs
Sales Salaries 150000
Advertising 253000
Others 98000
Fixed Administrative Costs
Administrative Staff Salaries 220000
Legal and Accounting Staff Salaries 80000
Data Processing services 65000
Outside professional services 52000
Depreciation 49000
Interest 6000
Property Tax 76000
Others 24000
Total Fixed Cost 1,557,000

Breakeven point in terms of units shall be - Total Fixed cost/contribution per unit

Contribution per unit = Selling price per unit - Estimated production cost per unit

= 40 - 32 = $8/unit.

Breakeven No. of units = 1557000/8 = 194625

2) Assuming that no. of units sold will be the same as in the above given situation i.e. 195000.

Hence, breakeven point would be calculated as -

Desired contribution per unit =Total Fixed Cost / No. of units

= 1557000/195000 = 7.9846

Breakeven selling price = Production cost + Desired contribution

= 32+ 7.9846

= 39.9846 Approx.

i.e. $40/unit if rounded up to nearest whole number.

3) Profit or loss when budgeted selling price is $ 30.00 and budgeted volume is $ 250000

Particulars Amount
Total Sales @ $30.00/unit 7,500,000
Production Cost @ 32/unit 8,000,000
Fixed Costs - as calculated in 1) above 1,557,000
Net profit (2,077,000)
Net profit after tax (2,077,000)

Since, there is net operating loss it would not be liable for any income taxes.

4) Profit or loss when budgeted selling price is $ 35.00 and budgeted volume is $ 150000

Particulars Amount
Total Sales @ $35.00/unit 5,250,000
Production Cost @ 32/unit (4,800,000)
Fixed Costs - as calculated in 1) above (1,557,000)
Net profit (1,107,000)
Net profit after tax (1,107,000)

Since, there is net operating loss it would not be liable for any income taxes.